# Parabolic Equations with Singular Coefficients and Boundary Data: Analysis and Numerical Simulations

### v1

@@ 2.2. Energy Estimates
$\qquad\times\left(\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|u_{0}\|_{L^{2}(0,1)}\right),$ (2.14)
$\Big\{\|\partial_{t}u\|_{L^{2}([0,T];L^{2})};\,\sqrt{2\alpha}\|\partial_{x}u\|_{L^{\infty}([0,T];L^{2})};\,\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2})};\,\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2})}\Big\}$
$\quad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q\|_{L^{\infty}(0,1)}$
$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\quad\times\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$ (2.15)
where the constant $C>0$ depends only on the final time $T$ and the spatial interval $(0,1)$, but not on the data or coefficients.
@@
By taking limits in (2.2)-(2.2) as $m=m_{l}\to\infty$, we derive the estimates (2.4)-(2.4).
To prove (2.4), we start from the equation in strong form:
$\partial_{t}u-\partial_{x}(a\partial_{x}u)=-b\partial_{x}u-qu+f.$
Multiply both sides by $\partial_{t}u$ and integrate over $(0,1)$:
$\|\partial_{t}u\|^{2}_{L^{2}}$ $-Re(\partial_{x}(a\partial_{x}u)\,,\partial_{t}u)+Re(b\partial_{x}u\,,\partial_{t}u)$
$+Re(qu\,,\partial_{t}u)=Re(f\,,\partial_{t}u).$
We handle the second term by integration by parts:
$-Re(\partial_{x}(a\partial_{x}u)\,,\partial_{t}u)=Re(a\partial_{x}u\,,\partial_{x}(\partial_{t}u))=\frac{1}{2}\frac{d}{dt}(a\,,|\partial_{x}u|^{2})-\frac{1}{2}(\partial_{t}a\,,|\partial_{x}u|^{2})$
$\qquad\geq\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}-\|\partial_{t}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)},$
where the second term we move to the right hand-side.
Apply Cauchy–Schwarz and Young’s inequality to the remaining terms:
$\left|(b\partial_{x}u\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|b\|^{2}_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}},$
$\left|(qu\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|q\|^{2}_{L^{\infty}}\|u\|^{2}_{L^{2}},$
$\left|(f\,,\partial_{t}u)\right|$ $\leq\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\|f\|^{2}_{L^{2}}.$
Combining everything, we obtain:
$\frac{1}{4}\|\partial_{t}u\|^{2}_{L^{2}}+\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u(t)\|^{2}_{L^{2}(0,1)}$ $\leq\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}}+\|b\|^{2}_{L^{\infty}}\|\partial_{x}u\|^{2}_{L^{2}}$
$+\|q\|^{2}_{L^{\infty}}\|u\|^{2}_{L^{2}}+\|f\|^{2}_{L^{2}}.$
Now integrate over $[0,T]$:
$\frac{1}{4}\|\partial_{t}u(t)\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u\|^{2}_{L^{\infty}([0,T];L^{2}(0,1))}$
$\qquad\leq\left(\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|^{2}_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\|\partial_{x}u\|^{2}_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad+\|q\|^{2}_{L^{\infty}(0,1)}\|u\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\|f\|^{2}_{L^{2}([0,T];L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u(0)\|^{2}_{L^{2}(0,1)}.$
We now estimate the last term $\|\partial_{x}u(0)\|_{L^{2}}$ by using initial condition, we obtain
$\|\partial_{x}u(0)\|_{L^{2}(0,1)}=\|\partial_{x}u_{0}\|_{L^{2}(0,1)}=\|u_{0}\|_{H^{1}_{0}(0,1)}.$
Using the full energy estimate, we also have bounds for $\|u\|_{L^{2}([0,T];L^{2})}$, $\|\partial_{x}u\|_{L^{2}([0,T];L^{2})}$ in terms of data. Therefore:
$\|\partial_{t}u\|_{L^{2}([0,T];L^{2}(0,1))}+\sqrt{2\alpha}\|\partial_{x}u\|_{L^{\infty}([0,T];L^{2}(0,1))}\leq C\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}$
$\qquad\times\left(\|a\|^{1/2}_{L^{\infty}}+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}}+\|b\|_{L^{\infty}}\right)+\|q\|_{L^{\infty}}\right)$
$\qquad\times\left(1+\frac{1}{\sqrt{\alpha}}\left(\|a\|_{L^{\infty}}+\|b\|_{L^{\infty}}\right)+\|q\|^{\frac{1}{2}}_{L^{\infty}}\right)\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}\right),$
where $C$ depends only on $T$ and the domain $(0,1)$.
Now, use the PDE again to write:
$\partial_{x}(a\partial_{x}u)=\partial_{t}u+b\partial_{x}u+qu-f.$
Then
$\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2})}$ $\leq\|\partial_{t}u\|_{L^{2}}+\|b\|_{L^{\infty}}\|\partial_{x}u\|_{L^{2}}+\|q\|_{L^{\infty}}\|u\|_{L^{2}}+\|f\|_{L^{2}},$
which again is controlled by data via the energy estimate.
Finally, to estimate the second spatial derivative $\partial_{x}^{2}u$, we use the identity
$a\,\partial_{x}^{2}u=\partial_{x}(a\,\partial_{x}u)-\partial_{x}a\,\partial_{x}u,$
and the ellipticity condition $a(t,x)\geq\alpha>0$. Then, by the triangle inequality and Hölder’s inequality:
$\|\partial_{x}^{2}u\|_{L^{2}([0,T];L^{2}(0,1))}$ $=\left\|\frac{1}{a}\left(\partial_{x}(a\,\partial_{x}u)-\partial_{x}a\,\partial_{x}u\right)\right\|_{L^{2}}$
$\leq\frac{1}{\alpha}\left(\|\partial_{x}(a\,\partial_{x}u)\|_{L^{2}([0,T];L^{2})}+\|\partial_{x}a\,\partial_{x}u\|_{L^{2}([0,T];L^{2})}\right)$
$\leq\frac{1}{\alpha}\left(\|\partial_{x}(a\,\partial_{x}u)\|_{L^{2}([0,T];L^{2})}+\|\partial_{x}a\|_{L^{\infty}([0,T]\times(0,1))}\|\partial_{x}u\|_{L^{2}([0,T];L^{2})}\right).$
Therefore, $\partial_{x}^{2}u\in L^{2}([0,T];L^{2}(0,1))$ with a constant depending only on $T$ and the interval $(0,1)$.
This completes the proof, with all terms estimated in terms of data and coefficients, and the constant depending only on $T$ and the interval $(0,1)$. ∎
3. Non-homogeneous Dirichlet condition
We now consider the parabolic equation in divergence form (1) with diffusion coefficient, drift term, and potential with initial condition (1.2) and non-homogeneous Dirichlet boundary conditions (1.3).
We decompose the solution as
$u(t,x)=w(t,x)+\psi(t,x),$ (3.1)
where $w$ satisfies homogeneous Dirichlet boundary conditions
$w(t,0)=w(t,1)=0,$
and $\psi$ is a lifting function defined by
$\psi(t,x):=(1-x)g_{0}(t)+xg_{1}(t),\quad(t,x)\in[0,T]\times(0,1).$
This function satisfies:
  • $\psi(t,0)=g_{0}(t),\quad\psi(t,1)=g_{1}(t)$,
  • $\partial_{x}\psi(t,x)=g_{1}(t)-g_{0}(t)$ is independent of $x$,
  • $\psi\in H^{1}([0,T];H^{1}(0,1))$ provided $g_{0},g_{1}\in H^{1}([0,T])$.
Substituting (3.1) into (1), we obtain the equation for $w$:
$\left\{\begin{array}[]{l}\partial_{t}w-\partial_{x}(a(t,x)\partial_{x}w)+b(t,x)\partial_{x}w+q(x)w=\tilde{f}(t,x),\\[2.0pt] w(0,x)=u_{0}(x)-\psi(0,x),\quad x\in(0,1),\\[2.0pt] w(t,0)=w(t,1)=0,\quad t\in[0,T],\end{array}\right$. (3.2)
where the inhomogeneous source term is given by
$\tilde{f}(t,x)=$ $f(t,x)-\partial_{t}\psi(t,x)+\partial_{x}(a(t,x)\partial_{x}\psi(t,x))-b(t,x)\partial_{x}\psi(t,x)-q(x)\psi(t,x)$
$=$ $f(t,x)-(1-x)g_{0}^{\prime}(t)-xg_{1}^{\prime}(t)+(\partial_{x}a(t,x)-b(t,x))(g_{1}(t)-g_{0}(t))$
$-q(x)\big((1-x)g_{0}(t)+xg_{1}(t)\big).$
The right-hand side $\tilde{f}\in L^{2}([0,T];L^{2}(0,1))$ under the assumptions $f\in L^{2}$ and $g_{0},g_{1}\in H^{1}([0,T])$, and the fact that the coefficients are bounded. Moreover, $\|\tilde{f}\|_{L^{2}}$ can be bounded by
$\|\tilde{f}\|_{L^{2}}\leq\|f\|_{L^{2}}+C\left(\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\right),$
for some constant $C>0$ depending only on $\|a\|_{W^{1,\infty}},\|b\|_{L^{\infty}},\|q\|_{L^{\infty}}$.
We now state the energy estimate.
Theorem 3.1.
Let $u$ solve the parabolic problem (1)–(1.3), with initial data $u_{0}\in L^{2}(0,1)$, source term $f\in L^{2}([0,T];L^{2}(0,1))$, and boundary data $g_{0},g_{1}\in H^{1}([0,T])$. Assume the coefficients satisfy
$a(t,x)\geq\alpha>0,\quad a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad q\in L^{\infty}(0,1).$
Then the solution $u\in L^{2}([0,T];H^{1}(0,1))$ with $\partial_{t}u\in L^{2}([0,T];H^{-1}(0,1))$ satisfies:
$\|u\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}u\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q}u\|_{L^{2}([0,T];L^{2}(0,1))}$
[…]
$\qquad\times\Big(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\Big),$ (3.3)
where $C>0$ depends only on $T$ and the domain $(0,1)$, but not on the data or coefficients.
Proof.
As described above, we write $u=w+\psi$, where $w$ satisfies the homogeneous problem (3.2). Applying the energy estimate for $w$ from Theorem 2.2, we obtain:
$\|w\|_{L^{\infty}L^{2}}+\sqrt{\alpha}\|\partial_{x}w\|_{L^{2}L^{2}}+\|\sqrt{q}w\|_{L^{2}L^{2}}\leq Ce^{T\|\partial_{x}b\|_{L^{\infty}}}\left(\|u_{0}-\psi(0)\|_{L^{2}}+\|\tilde{f}\|_{L^{2}}\right).$ (3.4)
Since $\psi(0,x)=(1-x)g_{0}(0)+xg_{1}(0)$, we estimate:
$\|\psi(0)\|_{L^{2}}\leq C\left(|g_{0}(0)|+|g_{1}(0)|\right)\leq C\left(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}\right),$
hence,
$\|u_{0}-\psi(0)\|_{L^{2}}\leq\|u_{0}\|_{L^{2}}+C(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}).$
Next, since $u=w+\psi$, we use the triangle inequality:
$\|u\|_{L^{\infty}L^{2}}$ $\leq\|w\|_{L^{\infty}L^{2}}+\|\psi\|_{L^{\infty}L^{2}},$
$\|\partial_{x}u\|_{L^{2}L^{2}}$ $\leq\|\partial_{x}w\|_{L^{2}L^{2}}+\|\partial_{x}\psi\|_{L^{2}L^{2}},$
$\|\sqrt{q}u\|_{L^{2}L^{2}}$ $\leq\|\sqrt{q}w\|_{L^{2}L^{2}}+\|\sqrt{q}\psi\|_{L^{2}L^{2}}.$
Note that $\psi(t,x)\in\text{span}\{1,x\}\cdot g_{j}(t)$, so
$\|\psi\|_{L^{\infty}L^{2}}+\|\partial_{x}\psi\|_{L^{2}L^{2}}+\|\sqrt{q}\psi\|_{L^{2}L^{2}}\leq C(T)\left(1+\|a\|^{1/2}_{L^{\infty}}+\|q\|^{1/2}_{L^{\infty}}\right)(\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}).$
Combining this with (3.4) yields (3.1), which completes the proof of Theorem 3.1. ∎
Corollary 3.2.
Let $u$ solve the non-homogeneous Dirichlet problem (1)–(1.3) with data
$f\in L^{2}([0,T];L^{2}(0,1)),\quad u_{0}\in L^{2}(0,1),\quad g_{0},g_{1}\in H^{1}([0,T]),$
and assume the coefficients satisfy:
$a,\,\partial_{t}a,\,\partial_{x}a,\,b,\,\partial_{x}b\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad a(t,x)\geq\alpha>0,\quad q\in L^{\infty}(0,1).$
Then:
$\partial_{t}u,\,\partial_{x}(a\partial_{x}u)\in L^{2}([0,T];L^{2}(0,1)),$
and the following estimate holds:
$\Big\{\|\partial_{t}u\|_{L^{2}([0,T];L^{2}(0,1))};\,\|\partial_{x}(a\partial_{x}u)\|_{L^{2}([0,T];L^{2}(0,1))};\,\|\partial^{2}_{x}u\|_{L^{2}([0,T];L^{2}(0,1))}\Big\}$
$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q\|_{L^{\infty}(0,1)}$
$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\qquad\times\left(\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0}\|_{H^{1}([0,T])}+\|g_{1}\|_{H^{1}([0,T])}\right),$
where the constant $C>0$ depends only on $T$ and the domain $(0,1)$, but not on the data or coefficients.
Proof.
We decompose the solution as $u=w+\psi$, where $w$ solves the corresponding homogeneous Dirichlet problem with modified source
$\tilde{f}=f-\mathcal{L}\psi,\qquad\psi(t,x)=(1-x)g_{0}(t)+xg_{1}(t)$
is the boundary lifting.
Then, since $\partial_{t}\psi\in L^{2}([0,T];L^{2}(0,1))$ and $\partial_{x}(a\partial_{x}\psi)\in L^{2}([0,T];L^{2}(0,1))$, and the estimate from Proposition 2.4 applies to $w$, we have:
$\|\partial_{t}u\|_{L^{2}}\leq\|\partial_{t}w\|_{L^{2}}+\|\partial_{t}\psi\|_{L^{2}},\quad\|\partial_{x}(a\partial_{x}u)\|_{L^{2}}\leq\|\partial_{x}(a\partial_{x}w)\|_{L^{2}}+\|\partial_{x}(a\partial_{x}\psi)\|_{L^{2}}.$
Now, note that $\psi\in H^{1}([0,T];H^{1}(0,1))$, and all terms involving $\psi$ can be bounded in terms of $\|g_{0}\|_{H^{1}}+\|g_{1}\|_{H^{1}}$. Combining with the bound for $w$ from Proposition 2.4 yields the desired result. ∎
4. Very weak solutions

### v2

@@ 2.2. Energy Estimates
Using these convergences and the boundedness of $a,b$ on $0,T]\times(0,1)$ and of $q$ on $(0,1)$, we can pass to the limit in (2.6) and conclude that $u$ satisfies (1). Moreover, by the standard Lions lemma (see [23, Lemma 1.2, p. 260]), $u\in L^{\infty}([0,T];L^{2}(0,1))$ and $u(0)=u_{0}$.
Uniqueness. Let $u$ be a weak solution with $u_{0}=0$ and $f=0$. Testing the weak formulation with $u(t)$ and taking real parts yields
$\operatorname{Re}\langle\partial_{t}u(t),u(t)\rangle+\operatorname{Re}(a\,\partial_{x}u(t),\partial_{x}u(t))+\operatorname{Re}(b\,\partial_{x}u(t),u(t))+\operatorname{Re}(qu(t),u(t))=0.$
Since $u\in L^{2}([0,T];H_{0}^{1})$ and $\partial_{t}u\in L^{2}([0,T];H^{-1})$, we have the identity
$\operatorname{Re}\langle\partial_{t}u(t),u(t)\rangle=\frac{1}{2}\frac{d}{dt}\|u(t)\|_{L^{2}(0,1)}^{2}\quad\text{in }\mathcal{D}^{\prime}(0,T).$
Moreover, integrating by parts gives
$\operatorname{Re}(b\,\partial_{x}u,u)=-\frac{1}{2}(\partial_{x}b,|u|^{2})\leq\frac{1}{2}\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|u\|_{L^{2}(0,1)}^{2}.$
Writing $q=q_{+}-q_{-}$ with $q_{+}\geq 0$ and $q_{-}:=\max\{-q,0\}$, we obtain
$\operatorname{Re}(qu,u)=(q_{+},|u|^{2})-(q_{-},|u|^{2})\geq\|\sqrt{q_{+}}\,u\|_{L^{2}(0,1)}^{2}-\|q_{-}\|_{L^{\infty}(0,1)}\|u\|_{L^{2}(0,1)}^{2}.$
Using also $\operatorname{Re}(a\partial_{x}u,\partial_{x}u)\geq\alpha\|\partial_{x}u\|_{L^{2}(0,1)}^{2}$, we arrive at
$\frac{1}{2}\frac{d}{dt}\|u(t)\|_{L^{2}(0,1)}^{2}+\alpha\|\partial_{x}u(t)\|_{L^{2}(0,1)}^{2}+\|\sqrt{q_{+}}\,u(t)\|_{L^{2}(0,1)}^{2}\leq C\,\|u(t)\|_{L^{2}(0,1)}^{2},$
where $C:=\tfrac{1}{2}\|\partial_{x}b\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|q_{-}\|_{L^{\infty}(0,1)}$. Since $u(0)=0$, Grönwall’s inequality implies $u\equiv 0$ on $[0,T]$. ∎
Lemma 2.4 (Higher time–space regularity).
Let $u$ be the unique weak solution of (2.1) given by Theorem 2.3 with initial data $u_{0}\in H^{1}_{0}(0,1)$ and $f\in L^{2}(0,T;L^{2}(0,1))$. Assume that the coefficients satisfy
$a,\,\partial_{t}a,\,\partial_{x}a,\,b,\,\partial_{x}b\in L^{\infty}(0,T;L^{\infty}(0,1)),\qquad a(t,x)\geq\alpha>0,$
and
$q\in L^{\infty}(0,1),\qquad q=q_{+}-q_{-},\quad q_{+}\geq 0,\quad q_{-}:=\max\{-q,0\}.$
Then the solution $u$ satisfies:
$\|u\|_{L^{\infty}(0,T;L^{2}(0,1))}+\sqrt{\alpha}\,\|\partial_{x}u\|_{L^{2}(0,T;L^{2}(0,1))}+\|\sqrt{q_{+}}\,u\|_{L^{2}(0,T;L^{2}(0,1))}$
$\qquad\leq C\exp\!\left\{T\big(\|\partial_{x}b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}$
$\qquad\quad\Big(\|u_{0}\|_{L^{2}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\Big),$ (2.14)
$\|\partial_{t}u\|_{L^{2}(0,T;H^{-1}(0,1))}\leq C\exp\!\left\{T\big(\|\partial_{x}b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}$
$\qquad\times\left(1+\frac{1}{\sqrt{\alpha}}\Big(\|a\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\Big)+\|q\|_{L^{\infty}(0,1)}\right)$
$\qquad\times\Big(\|f\|_{L^{2}(0,T;L^{2}(0,1))}+\|u_{0}\|_{L^{2}(0,1)}\Big),$ (2.15)
and moreover
$\|\partial_{t}u\|_{L^{2}(0,T;L^{2}(0,1))}+\sqrt{\alpha}\,\|\partial_{x}u\|_{L^{\infty}(0,T;L^{2}(0,1))}+\|\partial_{x}(a\partial_{x}u)\|_{L^{2}(0,T;L^{2}(0,1))}$
$\quad+\|\partial_{x}^{2}u\|_{L^{2}(0,T;L^{2}(0,1))}\leq\frac{C}{\alpha}\exp\!\left\{T\big(\|\partial_{x}b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\right\}$
$\qquad\times\Big(1+\|q\|_{L^{\infty}(0,1)}+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a\|_{L^{\infty}(0,T;L^{\infty}(0,1))}$
$\qquad+\|b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)\times\Big(\sqrt{\alpha}\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\Big),$ (2.16)
where $C>0$ depends only on $T$ and the interval $(0,1)$.
@@
The estimates (2.4)–(2.4) follow by passing to the limit in the Galerkin estimates (2.2).
We prove (2.4) at the Galerkin level and then pass to the limit. Let $u_{m}$ be the Galerkin approximations. Since $u_{m}\in C^{1}([0,T];V_{m})$, all computations below are justified. From the Galerkin equation we have (in the finite-dimensional sense)
$\partial_{t}u_{m}-\partial_{x}(a\partial_{x}u_{m})=-b\,\partial_{x}u_{m}-q\,u_{m}+f.$
Multiply by $\partial_{t}u_{m}$ and integrate over $(0,1)$:
$\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}$- $\operatorname{Re}\big(\partial_{x}(a\partial_{x}u_{m}),\partial_{t}u_{m}\big)$
$=-\operatorname{Re}\big(b\,\partial_{x}u_{m},\partial_{t}u_{m}\big)-\operatorname{Re}\big(q\,u_{m},\partial_{t}u_{m}\big)+\operatorname{Re}\big(f,\partial_{t}u_{m}\big).$
Integrating by parts,
$-\operatorname{Re}\big(\partial_{x}(a\partial_{x}u_{m}),\partial_{t}u_{m}\big)$ $=\operatorname{Re}\big(a\partial_{x}u_{m},\partial_{x}\partial_{t}u_{m}\big)$
$=\frac{1}{2}\frac{d}{dt}\,(a,|\partial_{x}u_{m}|^{2})-\frac{1}{2}(\partial_{t}a,|\partial_{x}u_{m}|^{2})$
$\geq\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}^{2}-\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\,\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}^{2}.$
Using Cauchy–Schwarz and Young’s inequality,
$\big|\big(b\partial_{x}u_{m},\partial_{t}u_{m}\big)\big|$ $\leq\frac{1}{4}\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}+\|b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}^{2}\|\partial_{x}u_{m}\|_{L^{2}(0,1)}^{2},$
$\big|\big(qu_{m},\partial_{t}u_{m}\big)\big|$ $\leq\frac{1}{4}\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}+\|q\|_{L^{\infty}(0,1)}^{2}\|u_{m}\|_{L^{2}(0,1)}^{2},$
$\big|\big(f,\partial_{t}u_{m}\big)\big|$ $\leq\frac{1}{4}\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}+\|f\|_{L^{2}(0,1)}^{2}.$
Combining these estimates yields
$\frac{1}{4}\|\partial_{t}u_{m}\|_{L^{2}(0,1)}^{2}+\frac{\alpha}{2}\frac{d}{dt}\|\partial_{x}u_{m}(t)\|_{L^{2}(0,1)}^{2}$ $\leq\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}}\|\partial_{x}u_{m}\|_{L^{2}(0,1)}^{2}+\|b\|_{L^{\infty}}^{2}\|\partial_{x}u_{m}\|_{L^{2}(0,1)}^{2}$
$\quad+\|q\|_{L^{\infty}}^{2}\|u_{m}\|_{L^{2}(0,1)}^{2}+\|f\|_{L^{2}(0,1)}^{2}.$
Integrating the above inequality over $(0,T)$ and taking the supremum with respect to $t\in[0,T]$, we obtain
$\frac{1}{4}\|\partial_{t}u_{m}\|^{2}_{L^{2}(0,T;L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u_{m}\|^{2}_{L^{\infty}(0,T;L^{2}(0,1))}$
$\qquad\leq\left(\frac{1}{2}\|\partial_{t}a\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b\|^{2}_{L^{\infty}(0,T;L^{\infty}(0,1))}\right)\|\partial_{x}u_{m}\|^{2}_{L^{2}(0,T;L^{2}(0,1))}$
$\qquad\quad+\|q\|^{2}_{L^{\infty}(0,1)}\|u_{m}\|^{2}_{L^{2}(0,T;L^{2}(0,1))}+\|f\|^{2}_{L^{2}(0,T;L^{2}(0,1))}+\frac{\alpha}{2}\|\partial_{x}u_{m}(0)\|^{2}_{L^{2}(0,1)}.$
Using the initial condition $u_{m}(0)=P_{m}u_{0}$ and the stability of the projection $P_{m}$ in $H^{1}_{0}(0,1)$, we have
$\|\partial_{x}u_{m}(0)\|_{L^{2}(0,1)}\leq\|u_{0}\|_{H^{1}_{0}(0,1)}.$
Moreover, the energy estimate (2.4) yields bounds for $\|u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}$ and
$\|\partial_{x}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}$ in terms of the data. Consequently, there exists a constant $C>0$, independent of $m$, such that
$\|\partial_{t}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}+\sqrt{\alpha}\,\|\partial_{x}u_{m}\|_{L^{\infty}(0,T;L^{2}(0,1))}$
$\quad\leq C\exp\!\Big\{T\big(\|\partial_{x}b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-}\|_{L^{\infty}(0,1)}\big)\Big\}$
$\qquad\times\Big(1+\|q\|_{L^{\infty}(0,1)}+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
$\qquad\times\Big(\sqrt{\alpha}\,\|u_{0}\|_{H^{1}_{0}(0,1)}+\|f\|_{L^{2}(0,T;L^{2}(0,1))}\Big),$ (2.17)
where $C$ depends only on $T$.
The previous estimate is uniform in $m$, hence $\{\partial_{t}u_{m}\}$ is bounded in $L^{2}(0,T;L^{2}(0,1))$. By Banach–Alaoglu, there exist a subsequence (not relabelled) and a function $w\in L^{2}(0,T;L^{2}(0,1))$ such that
$\partial_{t}u_{m}\rightharpoonup w\qquad\text{in }L^{2}(0,T;L^{2}(0,1)).$
On the other hand, from the energy estimate (2.4) we already know that $\{\partial_{t}u_{m}\}$ is bounded in $L^{2}(0,T;H^{-1}(0,1))$ and, along the same subsequence,
$\partial_{t}u_{m}\rightharpoonup\partial_{t}u\qquad\text{in }L^{2}(0,T;H^{-1}(0,1)).$
Since the embedding $L^{2}(0,1)\hookrightarrow H^{-1}(0,1)$ is continuous, the weak limit in
$L^{2}(0,T;L^{2}(0,1))$ is also the weak limit in $L^{2}(0,T;H^{-1}(0,1))$. Therefore $w=\partial_{t}u$. In particular,
$\partial_{t}u\in L^{2}(0,T;L^{2}(0,1))\quad\text{and}\quad\|\partial_{t}u\|_{L^{2}(0,T;L^{2}(0,1))}\leq\liminf_{m\to\infty}\|\partial_{t}u_{m}\|_{L^{2}(0,T;L^{2}(0,1))}.$
Similarly, by weak-* compactness in $L^{\infty}(0,T;L^{2}(0,1))$ we have (up to a subsequence)
$\partial_{x}u_{m}\stackrel{{\scriptstyle*}}{{\rightharpoonup}}\partial_{x}u\qquad\text{in }L^{\infty}(0,T;L^{2}(0,1)),$
and hence
$\|\partial_{x}u\|_{L^{\infty}(0,T;L^{2}(0,1))}\leq\liminf_{m\to\infty}\|\partial_{x}u_{m}\|_{L^{\infty}(0,T;L^{2}(0,1))},$
by weak-* lower semicontinuity of the $L^{\infty}$-norm. Consequently, the first two terms in (2.4) pass to the limit and the same bound holds for $u$.
Next, from the equation we write
$\partial_{x}(a\partial_{x}u)=\partial_{t}u+b\,\partial_{x}u+q\,u-f,$
so
$\|\partial_{x}(a\partial_{x}u)\|_{L^{2}(0,T;L^{2})}\leq$ $\|\partial_{t}u\|_{L^{2}(0,T;L^{2})}+\|b\|_{L^{\infty}}\|\partial_{x}u\|_{L^{2}(0,T;L^{2})}$
$+\|q\|_{L^{\infty}}\|u\|_{L^{2}(0,T;L^{2})}+\|f\|_{L^{2}(0,T;L^{2})},$
which is controlled by the previous bounds and (2.4). Finally,
$a\,\partial_{x}^{2}u=\partial_{x}(a\partial_{x}u)-(\partial_{x}a)\,\partial_{x}u,$
hence using $a\geq\alpha$,
$\|\partial_{x}^{2}u\|_{L^{2}(0,T;L^{2})}\leq\frac{1}{\alpha}\Big(\|\partial_{x}(a\partial_{x}u)\|_{L^{2}(0,T;L^{2})}+\|\partial_{x}a\|_{L^{\infty}}\|\partial_{x}u\|_{L^{2}(0,T;L^{2})}\Big),$
and (2.4) follows by collecting the estimates and absorbing numerical factors into $C$. ∎
3. Very weak solutions for the non-homogeneous Dirichlet conditions